# Controllable Generation with Text-to-Image Diffusion Models: A Survey

### v1

@@ Abstract
In the rapidly advancing realm of visual generation, diffusion models have revolutionized the landscape, marking a significant shift in capabilities with their impressive text-guided generative functions. However, relying solely on text for conditioning these models does not fully cater to the varied and complex requirements of different applications and scenarios. Acknowledging this shortfall, a variety of studies aim to control pre-trained text-to-image (T2I) models to support novel conditions. In this survey, we undertake a thorough review of the literature on controllable generation with T2I diffusion models, covering both the theoretical foundations and practical advancements in this domain. Our review begins with a brief introduction to the basics of denoising diffusion probabilistic models (DDPMs) and widely used T2I diffusion models. We then reveal the controlling mechanisms of diffusion models, theoretically analyzing how novel conditions are introduced into the denoising process for conditional generation. Additionally, we offer a detailed overview of research in this area, organizing it into distinct categories from the condition perspective: generation with specific conditions, generation with multiple conditions, and universal controllable generation. For an exhaustive list of the controllable generation literature surveyed, please refer to our curated repository at https://github.com/PRIV-Creation/Awesome-Controllable-T2I-Diffusion-Models.
@@ 1 Introduction
Diffusion models, representing a paradigm shift in the visual generation, have dramatically outperformed traditional frameworks like Generative Adversarial Networks (GANs)[1, 2, 3, 4, 5, 6, 7, 8]. As parameterized Markov chains, diffusion models exhibit a remarkable ability to transform random noise into intricate images, progressing sequentially from noise to high-fidelity visual representations. With the advancement of technology, diffusion models have demonstrated immense potential in image generation and related downstream tasks.
As the quality of imagery generated by these models advances, a critical challenge becomes increasingly apparent: achieving precise control over these generative models to fulfill complex and diverse human needs. This task goes beyond simply enhancing image resolution or realism; it involves meticulously aligning the generated output with the user’s specific and nuanced requirements as well as their creative aspirations. Fueled by the advent of extensive multi-modal text-image datasets [9, 10, 11, 12, 13, 14, 15, 16, 17] and development of guidance mechanism [18, 19, 20, 21], text-to-image (T2I) diffusion models have emerged as a cornerstone in the controllable visual generation landscape [22, 21, 23, 24, 25, 26]. These models are capable of generating realistic, high-quality images that accurately reflect the descriptions provided in natural language.
(a) Half-yearly paper count.
@@
While text-based conditions have been instrumental in propelling the field of controllable generation forward, they inherently lack the capability to fully satisfy all user requirements. This limitation is particularly evident in scenarios where conditions, such as the depiction of an unseen person or a distinct art style, are not effectively conveyable through text prompts alone. These scenarios pose significant challenges in the T2I generation process, as the nuances and complexities of such visual representations are difficult to encapsulate in text form. Recognizing this gap, a substantial body of research has shifted focus towards integrating novel conditions that extend beyond the confines of textual descriptions into T2I diffusion models. This pivot has been further facilitated by the emergence of powerful and open-sourced T2I diffusion models, as illustrated in Figure 1a. These advancements have led to the exploration of diverse conditions, thereby enriching the spectrum of possibilities for conditional generation and addressing the more intricate and nuanced demands of users in various applications.
There are numerous survey articles exploring the AI-generated content (AIGC) domain, including diffusion model theories and architectures [28], efficient diffusion models [29], multi-modal image synthesis and editing [30], visual diffusion model [31, 32, 33, 34], and text-to-3D applications [35]. However, they often provide only a cursory brief of controlling text-to-image diffusion models or predominantly focus on alternative modalities. This lack of in-depth analysis of the integration and impact of novel conditions in T2I models highlights a critical area for future research and exploration.
This survey provides an exhaustive review of controllable generation using text-to-image diffusion models, encompassing both theoretical foundations and practical applications. Initially, we provide a concise overview of the background of T2I diffusion models and delve into the theoretical underpinnings of these methods, elucidating how novel conditions are integrated into T2I diffusion models. This exploration sheds light on the fundamentals of prior research and facilitates a deeper understanding of the field. Subsequently, we offer a thorough overview of previous studies, highlighting their unique contributions and distinguishing features. Additionally, we explore the varied applications of these methods, showcasing their practical utility and impact in diverse contexts and related tasks.
@@
  • We introduce a well-structured taxonomy of controllable generation methods from the condition perspective, shedding light on the inherent challenges and complexities in this study area.
  • We conduct an in-depth analysis of two core theoretical mechanisms essential for incorporating novel conditions into T2I diffusion models: conditional score prediction and condition-guided score estimation, providing a nuanced understanding of how these mechanisms function at a granular level.
  • Our review is comprehensive, covering a wide range of conditional generation studies according to our proposed taxonomy. We meticulously underscore the salient features and distinctive characteristics of each method.
  • We showcase the diverse applications of conditional generation using T2I diffusion models across various generative tasks, demonstrating its emergence as a fundamental and influential aspect in the AIGC era.
The rest of this paper is organized as follows. Section 2 provides a brief introduction to denoising diffusion probabilistic models (DDPMs), demonstrates the widely used text-to-image diffusion models, and presents a well-structured taxonomy. In Section 3, we analyze the controlling mechanisms and reveal how to introduce novel conditions in text-to-image diffusion models. In Section 4, we summarize existing approaches for controlling the text-to-image diffusion model according to our proposed taxonomy. Finally, section 7 demonstrates the applications of controllable text-to-image generation.
@@ 2.1 Denoising Diffusion Probabilistic Models
where deep neural networks, often instantiated by architectures like UNet, parameterize the mean $\mu_{\theta}(x_{t},t)$ and variance $\Sigma_{\theta}(x_{t},t)$. The UNet takes the noised data $x_{t}$ and time step $t$ as inputs and outputs the parameters of the normal distribution, thereby predicting the noise $\epsilon_{\theta}$ that the model needs to reverse the diffusion process. To synthesize new data instances $x_{0}$, we initiate by sampling a noise vector $x_{T}\sim p(x_{T})$ and then successively sample from the learned transition kernels $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t})$ until we reach $t=1$, completing the reverse diffusion process.
2.2 Text-to-Image Diffusion Models
In this section, we spotlight several pivotal and widely utilized text-to-image foundational models. Detailed information regarding these models is systematically compiled and presented in Table I.
@@ 2.2 Text-to-Image Diffusion Models
Table I: Collection of primary and used text-to-image diffusion models in this survey. ${}^{\dagger}$: number of UNet and text encoder’s parameters (default refers only to UNet). $f$: downsampling factor of autoencoder in latent-space diffusion models. CLIP: open source implementation of CLIP. $*{}^{*}$: train from scratch.
  Model Publication Param. Resolution
$f$
Text Encoder Training Dataset Open Source
Pixel Space Diffusion Models
GLIDE[21] ICML 2022 ${}^{\dagger}$ $256^{2}$
[…]
Imagen[24] NeurIPS 2022 3.0B $1024^{2}$
T5-XXL[37] $>$LAION-400M[16]
@@
LDM[23] CVPR 2022 903M $256^{2}$
$8$
BERT-tokenizer[40] LAION-400M[16]
SD v1.x[23] CVPR 2022 860M $512^{2}$
8
CLIP-ViT-L/14[39] LAION-2B[17]
SD v2.x[23] CVPR 2022 865M $512^{2}$/$768^{2}$
8
CLIP-ViT-H/14[39] LAION-5B[17]
SD XL[25] ICLR 2024 2.6B $1024^{2}$
8
CLIP-ViT/G & CLIP-ViT/L[39] internal dataset
• Imagen [24]. Following GLIDE, Imagen adopts classifier-free guidance (CFG) for text-to-image generation. Instead of training a task-specified text encoder from scratch in GLIDE, Imagen leverages a pre-trained and frozen large language model (LLM) as its text encoder, aiming to reduce computational demands. The authors conduct a comparative analysis of various LLMs, including those trained on image-text datasets (e.g., CLIP[39]) and solely on text corpora (e.g., BERT[40], T5[37]). Their findings suggest that increasing the scale of language models more effectively enhances the fidelity of samples and the congruence between image and text, compared to the enlargement of image diffusion models. Furthermore, Imagen’s exploration into different text conditioning methods reveals cross-attention as the most effective technique.
@@
• Latent Diffusion Model (LDM) [23]. To enable diffusion model training and inference on limited computational resources and generate high-resolution images of high quality and flexibility, LDM applies the denoising process in the latent space of pre-trained autoencoders. Specifically, the autoencoder $\mathcal{E}$ maps images $x\in\mathcal{D}_{x}$ into a spatial latent space $z=\mathcal{E}(x)$. To develop a conditional image generator, LDM enhances the underlying UNet with the cross-attention mechanism to effectively model the conditional distribution $p(z_{t-1}|z_{t},c)$, where $c$ is the conditional input, such as text prompts and segmentation masks.
In the realm of text-to-image generation, the authors employ the LAION-400M dataset to train a 1.45 billion parameter text-to-image LDM model, capable of producing images at a resolution of $256\times 256$ (with a latent resolution of $32\times 32$). For the encoding of text inputs, a BERT tokenizer[40] was utilized as the text encoder.
• Stable Diffusion (SD). Built upon the Latent Diffusion Model (LDM) framework, Stability AI developed and launched several series of text-to-image diffusion models, termed Stable Diffusion. SD demonstrates unparalleled capabilities in text-to-image generation, and with its models being open-sourced, it has gained widespread usage within the community.

### v2

@@ Abstract
In the rapidly advancing realm of visual generation, diffusion models have revolutionized the landscape, marking a significant shift in capabilities with their impressive text-guided generative functions. However, relying solely on text for conditioning these models does not fully cater to the varied and complex requirements of different applications and scenarios. Acknowledging this shortfall, a variety of studies aim to control pre-trained text-to-image (T2I) models to support novel conditions. In this survey, we undertake a thorough review of the literature on controllable generation with T2I diffusion models, covering both the theoretical foundations and practical advancements in this domain. Our review begins with a brief introduction to the basics of denoising diffusion probabilistic models (DDPMs) and widely used T2I diffusion models. Additionally, we provide a detailed overview of research in this area, categorizing it from the condition perspective into three directions: generation with specific conditions, generation with multiple conditions, and universal controllable generation. For each category, we analyze the underlying control mechanisms and review representative methods based on their core techniques. For an exhaustive list of the controllable generation literature surveyed, please refer to our curated repository at https://github.com/PRIV-Creation/Awesome-Controllable-T2I-Diffusion-Models.
@@ 1 Introduction
Diffusion models, representing a paradigm shift in the visual generation, have dramatically outperformed traditional frameworks like Generative Adversarial Networks (GANs)[74, 120, 121, 119]. As parameterized Markov chains, diffusion models exhibit a remarkable ability to transform random noise into intricate images, progressing sequentially from noise to high-fidelity visual representations. With the advancement of technology, diffusion models have demonstrated immense potential in image generation and related downstream tasks.
As the quality of imagery generated by these models advances, a critical challenge becomes increasingly apparent: achieving precise control over these generative models to fulfill complex and diverse human needs. This task goes beyond simply enhancing image resolution or realism; it involves meticulously aligning the generated output with the user’s specific and nuanced requirements as well as their creative aspirations. Fueled by the advent of extensive multi-modal text-image datasets [128, 149, 229, 228] and development of guidance mechanism [57, 242, 93, 184], text-to-image (T2I) diffusion models have emerged as a cornerstone in the controllable visual generation landscape [214, 184, 219, 225, 199, 13]. These models are capable of generating realistic, high-quality images that accurately reflect the descriptions provided in natural language.
(a) Yearly paper count.
@@
While text-based conditions have been instrumental in propelling the field of controllable generation forward, they inherently lack the capability to fully satisfy all user requirements. This limitation is particularly evident in scenarios where conditions, such as the depiction of an unseen person or a distinct art style, are not effectively conveyable through text prompts alone. These scenarios pose significant challenges in the T2I generation process, as the nuances and complexities of such visual representations are difficult to encapsulate in text form. Recognizing this gap, a substantial body of research has shifted focus towards integrating novel conditions that extend beyond the confines of textual descriptions into T2I diffusion models. This pivot has been further facilitated by the emergence of powerful and open-sourced T2I diffusion models, as illustrated in Fig. 1a. These advancements have led to the exploration of diverse conditions, thereby enriching the spectrum of possibilities for conditional generation and addressing the more intricate and nuanced demands of users in various applications.
There are numerous survey articles exploring the AI-generated content (AIGC) domain, including diffusion model theories and architectures [303], efficient diffusion models [256], multi-modal image synthesis and editing [315], visual diffusion model [51, 297], and text-to-3D applications [135]. However, they often provide only a cursory brief of controlling text-to-image diffusion models or predominantly focus on alternative modalities. This lack of in-depth analysis of the integration and impact of novel conditions in T2I models highlights a critical area for future research and exploration.
This survey presents a comprehensive review of controllable generation with text-to-image diffusion models, covering both theoretical foundations and practical advancements. We begin with a concise overview of T2I diffusion models, introducing a brief summary of the theory and widely adopted text-to-image models. We then provide an in-depth examination of prior studies from a technical perspective, analyzing their theoretical underpinnings and highlighting their distinctive contributions and characteristics. This discussion not only clarifies the foundations of earlier research but also deepens the understanding of the field. Furthermore, we review the diverse applications of these methods, demonstrating their practical value and influence across different contexts and related tasks.
@@
  • We introduce a well-structured taxonomy of controllable generation methods from the condition perspective, encompassing novel condition introduction, multi-condition integration, and universal controllable generation. This taxonomy provides a clearer lens to reveal their core theoretical principles as well as the inherent challenges of each category.
  • We summarize two fundamental paradigms for incorporating novel conditions into T2I diffusion models: conditional score prediction and condition-guided score estimation. Based on these paradigms, we systematically organize and review the corresponding methods, covering a broad spectrum of controllable generation studies. We carefully highlight the key features, distinctive contributions, and comparative advantages of each method.
  • We further showcase the diverse applications of conditional generation with T2I diffusion models across a variety of generative tasks, illustrating its emergence as a central and influential component in the AIGC era.
The rest of this paper is organized as follows. Sec. 2 provides a brief introduction to the theory and widely used text-to-image diffusion models, overviews the popular controllable generation tasks, and presents a well-structured taxonomy. Later, we summarize existing approaches for controlling the text-to-image diffusion model according to our proposed taxonomy, including novel condition introduction (Sec. 3), multi-condition integration (Sec. 4), and universal controllable generation (Sec. 5). Finally, Sec. 6 demonstrates the applications of controllable text-to-image generation.
@@ 2.1 Denoising Diffusion Probabilistic Models
where deep neural networks, often instantiated by architectures like UNet, parameterize the mean $\mu_{\theta}(x_{t},t)$ and variance $\Sigma_{\theta}(x_{t},t)$. The UNet takes the noisy data $x_{t}$ and time step $t$ as inputs and outputs the parameters of the normal distribution, thereby predicting the noise $\epsilon_{\theta}$ that the model needs to reverse the diffusion process. To synthesize new data instances $x_{0}$, we initiate by sampling a noise vector $x_{T}\sim p(x_{T})$ and then successively sample from the learned transition kernels $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t})$ until we reach $t=1$, completing the reverse diffusion process.
Subsequent research has extended the DDPM framework toward more general and theoretically grounded formulations, including the Denoising Diffusion Implicit Model (DDIM) [240], score-based generative modeling [242], Flow Matching [152], and Elucidated Diffusion Models (EDM) [118]. Among these, we focus on the flow-matching technique, as it underpins many state-of-the-art text-to-image systems such as Stable Diffusion 3 [61] and Flux [132].
2.2 Flow Matching and Rectified Flow
• Flow Matching (FM) [152]. Flow Matching provides a deterministic reformulation of diffusion-based generative modeling by directly learning a continuous vector field that transports probability mass from the data distribution to a simple prior. Let $x_{t}$ denote the data state at time $t\in[0,1]$ evolving under
$\frac{dx_{t}}{dt}=u_{t}(x_{t}),\quad x_{0}\sim p_{\text{data}},\quad x_{1}\sim p_{1},$ (4)
where $v_{t}(x)$ represents the velocity field. To enable supervised training, a conditional Gaussian path $p_{t}(x\mid x_{0})$ is defined with analytical velocity
$u_{t}(x_{t}\mid x_{0})=\frac{\dot{\sigma}_{t}}{\sigma_{t}}\big(x_{t}-\mu_{t}\big)+\dot{\mu}_{t},$ (5)
parameterized by time-dependent schedules $\mu_{t}$ and $\sigma_{t}$ (which may depend on $x_{0}$). A neural network $u_{\theta}(x_{t},t)$ is trained to approximate this target velocity via the flow matching loss:
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{x_{0},\epsilon,t}\!\left[\|u_{\theta}(x_{t},t)-u_{t}(x_{t}|x_{0})\|^{2}\right],$ (6)
where $x_{t}=\mu_{t}(x_{0})+\sigma_{t}\epsilon$. Here, $\mu_{t}(x_{0})$ denotes a deterministic interpolation function, distinct from the learnable mean $\mu_{\theta}(x_{t},t)$ used in DDPMs. After training, generation proceeds deterministically by integrating the learned ODE from noise to data, removing stochasticity and improving sampling efficiency compared with DDPMs.
• Rectified Flow (RF) [161]. Rectified Flow simplifies the FM formulation by adopting a straight-line interpolation between the prior and the data:
$x_{t}=(1-t)x_{1}+tx_{0},\qquad u_{t}(x_{t}\mid x_{0})=x_{0}-x_{1},$ (7)
with $x_{0}\sim p_{\text{data}}$, $x_{1}\sim p_{1}$, and $t\sim\mathcal{U}(0,1)$. The training objective remains a mean-squared error on the predicted velocity:
$\mathcal{L}_{\mathrm{RF}}=\mathbb{E}_{x_{0},x_{1},t}\!\left[\|\,u_{\theta}(x_{t},t)-(x_{0}-x_{1})\|^{2}\right].$ (8)
This rectification enforces a monotonic and low-curvature probability flow, resulting in smoother vector fields and faster ODE integration.
Table I: Collection of primary and used text-to-image diffusion models in this survey. : number of UNet and text encoder’s parameters (default refers only to UNet). $f$: downsampling factor of autoencoder in latent-space diffusion models. CLIP: open source implementation of CLIP. : train from scratch. Resolution: the maximum supported image resolution generated by the model.
  Model Pub. Param. Resolution $f$ Text Encoder Arch. Training Dataset Open
Pixel Space Diffusion Models
@@ 2.2 Flow Matching and Rectified Flow
Imagen[225] NeurIPS 2022 3.0B $1024^{2}$ - T5-XXL[209] U-Net >>LAION-400M[229]
@@
LDM[219] CVPR 2022 903M $256^{2}$ 88 BERT-tokenizer[56] U-Net LAION-400M[229]
SD 1.x[219] CVPR 2022 860M $512^{2}$ 8 CLIP-L[208] U-Net LAION-2B[228]
SD 2.x[219] CVPR 2022 865M $512^{2}$$768^{2}$ 8 CLIP-H/14[208] U-Net LAION-5B[228]
SD XL[199] ICLR 2024 2.6B $1024^{2}$ 8 CLIP-G & CLIP-L[208] U-Net internal dataset
@@
PixArt-$\alpha$[35] ICLR 2024 0.25B $1024^{2}$ 8 T5-XXL[209] DiT mixture
Flux.1x [132] arXiv 2025 12B $2048^{2}$ 8 CLIP-L[208] & T5-XXL[209] $M$-DiT internal dataset
2.3 Text-to-Image Diffusion Models
Text-to-image (T2I) diffusion models synthesize images from textual descriptions by learning a conditional denoising process. A central design question in such models is how textual information is injected into the denoising network. We highlight here the representative architectural paradigms that define modern T2I systems regarding task formulation and architecture-level analysis (Tab. I).
• Cross-Attention–Based U-Net Architectures. Most early and widely adopted T2I models employ a U-Net backbone augmented with cross-attention layers to model. In this mechanism, latent image features act as queries, while text embeddings provide keys and values, enabling fine-grained alignment between linguistic concepts and visual features. GLIDE [184] first demonstrated that replacing class-conditioning in diffusion models with free-form text, combined with classifier-free guidance (CFG) [93], yields strong improvements in photorealism and text alignment. Imagen [225] further showed that scaling the text encoder (a frozen large language model, e.g., T5 [209]) improves generation quality more than enlarging the diffusion model itself, and confirmed cross-attention as the most effective conditioning approach. Latent Diffusion Models (LDM) [219] introduced a major efficiency breakthrough by performing diffusion in a compressed latent space, enabling high-resolution synthesis on limited compute. Stable Diffusion (SD) [219] builds on the LDM formulation and its v1.x, v2.x, and SDXL variants adopt the U-Net with cross-attention design that has become the de facto standard in open-source T2I generation.
• Transformer-Based Diffusion Models: DiT and MMDiT. Recent T2I systems, such as Stable Diffusion 3 [61] and FLUX [132], increasingly replace U-Nets with transformer-based diffusion backbones due to their improved scalability and expressiveness. Diffusion Transformers (DiT) [193] treat image latents as patch tokens and perform denoising entirely through self-attention, showing superior scaling behavior compared to convolutional models. Beyond traditional cross-attention, Multimodal Diffusion Transformers (MMDiT) introduce joint attention mechanisms in which text and image tokens interact bidirectionally within the same transformer blocks. This unified multimodal modeling enables richer text–image dependency structures than unidirectional cross-attention.
2.4 Controllable Generation Tasks
Figure 2: Illustration of controllable text-to-image generation with specific conditions. The condition is marked in blue background. Examples are sourced from [283, 37, 29, 103, 212, 22, 323, 277, 285, 167, 205, 322, 17, 251].
Controllable generation extends the capabilities of text-to-image diffusion models by introducing diverse conditions that guide the synthesis process beyond plain textual prompts. These conditions enable more precise alignment with user intentions. In this part, we introduce the major categories of controllable generation tasks and illustrate them in Fig. 2:
• Spatial Control. Since text alone struggles to represent structural information such as position and dense labels, spatial signals have become crucial conditions for text-to-image diffusion. Typical spatial inputs include layout, human pose, depth, and segmentation masks.
• Image Personalization. The personalization task aims to capture and utilize concepts from exemplar images that cannot be easily described by text, integrating them as generative conditions for controllable synthesis.
• View-conditioend Generation. View-conditioned generation aims to synthesize images from specific viewpoints or across multiple views, ensuring geometric consistency and structural coherence. By leveraging conditions such as camera parameters, depth maps, or multi-view correspondences, these methods extend controllable diffusion to scenarios like novel view synthesis, 3D-aware image generation, and panoramic rendering.
• Advanced Text-Conditioned Generation. Although text is the fundamental condition in text-to-image diffusion models, several challenges remain. For example, text-guided synthesis often suffers from misalignment, particularly when dealing with complex prompts involving multiple entities or rich contextual descriptions. Moreover, the dominance of English datasets in training leads to limited multilingual capabilities. To address these issues, recent works propose novel strategies for improving alignment and expanding the linguistic scope of text-conditioned generation.
• In-Context Generation. In-context generation focuses on understanding and executing specific tasks on query images by leveraging task-specific example pairs of images and text guidance. This setting enables models to adapt to new tasks with minimal supervision, expanding the flexibility of controllable diffusion.
• Brain- & Sound-guided Generation. Brain-guided generation seeks to control image synthesis directly from neural activity, such as electroencephalogram (EEG) recordings or functional magnetic resonance imaging (fMRI), bypassing the intermediate step of translating thoughts into text. Similarly, sound-guided generation explores how auditory signals can be leveraged as direct conditions for visual synthesis, enabling cross-modal creativity.
• Text Rendering. Rendering coherent and legible text within generated images is a critical task, given its wide applications in posters, book covers, advertisements, and memes. Effective text rendering not only enhances practicality but also pushes the boundary of fine-grained controllability in diffusion models.
Table II: Summary of evaluation dimensions and commonly used metrics for controllable T2I models.
Dimension Metrics
Image Quality FID [92]; IS [226]; Aesthetic Score[133];Human Preference(HPS [286],PickScore [125])
Conditional Alignment Textual Similarity(CLIP [91, 208]); Visual Similarity(CLIPScore[208], DINO [185]); Spatial Consistency(mIoU, MSE); Content Accuracy[100, 72]
2.5 Evaluation of Controllable T2I Generation
Since the types of conditions used in controllable T2I methods are highly diverse and vary across approaches, we focus here on the overarching evaluation dimensions rather than specific task settings. In what follows, we outline the main perspectives from which controllable T2I models are commonly assessed and summarize the widely adopted metrics associated with each dimension (Tab. II).
• Image Quality. Image quality is commonly measured by distribution-based metrics such as Inception Score (IS) [226] and Fréchet Inception Distance (FID) [92], computed on deep features from a pretrained classifier (e.g., Inception V3 [249]). In addition, aesthetic predictors [133] and learned human-preference models (HPS [286], PickScore [125]) estimate visual appeal and perceived quality directly from generated images.
• Conditional Alignment. Conditional alignment measures how faithfully the generated images follow the specified conditions (textual or otherwise). Text–image alignment is typically quantified with CLIP-based similarity scores [91, 208], which leverage a large-scale contrastively trained vision–language model to compare prompts and outputs. For other types of conditions, alignment is computed as the discrepancy between the input condition and its prediction from the generated image, for example by using CLIP or DINOv2 [185] features for image-conditioned tasks, or detection metrics such as mAP for layout-conditioned generation.